# A Review of Large Language Models and Autonomous Agents in Chemistry

### v1

@@ 3.6 Textual scientific LLMs
LLMs are large deep neural network models. Therefore, their superior performance against other smaller models in classical machine learning tasks is somewhat expected. Addressing tasks that seemed impossible to tackle in the past better illustrates LLM’s outstanding capabilities. That is, despite accurately predicting properties given well-structured information such as molecular features and descriptors, LLMs can access less structured data and extract information from simple natural language. For instance, instead of developing a restricted representation, Huang et al. 166 used clinical annotations from the MIMIC-III313 dataset to predict patient readmission. ClinicalBERT implements a pretrained BERT architecture using a combination of masked language modeling and next-sentence prediction training schemes followed by supervised fine-tuning on readmission prediction. Or, instead of pretraining, Zhao et al. 150 fine-tuned LLaMA on epilepsy data using an instruction-following approach to create EpilepsyLLM. Similarly, SciBERT162 and ScholarBERT165 adapted BERT to scientific literature. Beltagy et al. 162 created a tokenizer specific for scientific texts extracted from scientific articles from semantic scholar134 and showed that SciBERT outperforms fine-tuned BERT-base98 in every task investigated. It shows that vocabulary quality plays an important role in the model’s performance. A few years later, Hong et al. 165 pretrained BERT using RoBERTa314 optimizations to improve pretraining performance. ScholarBERT was pretrained on scientific articles from Public.Resource.Org, Inc and further fine-tuned on the tasks used for evaluation. Despite using a much larger dataset, ScholarBERT did not outperform other LLMs trained with narrower domain datasets. Even against SciBERT, another LLM trained on multi-domain articles, ScholarBERT could not perform relatively better in tasks from biomedical, computer science, and material domains. However, ScholarBERT showed significant improvement in the ScienceExamCER315 dataset, which contains semantic class annotations in a training corpus of 3rd to 9th grade science exam questions for named entity recognition (NER) tasks.
@@
On the other hand, instead of pretraining LLMs on a large domain scope, models focused on only one field were also developed. Guo et al. 169 argue that manually curating structured datasets is a sub-optimal, time-consuming, and labor-intensive task. Therefore, they automated data extraction and annotation from scientific papers using ChemDataExtractor316 and their in-house annotation tool317. ChemBERT169 was pretrained using a BERT model to encode chemical reaction information, followed by fine-tuning a NER head. ChemBERT outperformed other models such as BERT98 and BioBERT171 in the product extraction task, presenting an improvement of $\sim$ 6% in precision. For the product role labeling task – given the extracted compound, which role does it play in the reaction? –, it increases precision in $\sim$ 5% when compared to the same models. The increase in performance might be due to training in a narrower domain, which allows the model to learn specific traits from the data instead of being more generalist. MatSciBERT160 and MaterialsBERT167 followed a similar strategy. Gupta et al. 160 developed MatSciBERT by fine-tuning SciBERT162 on the Material Science Corpus (MSC), their curated dataset focused on materials extracted from Elsevier’s scientific papers. MatSciBERT shows an accuracy improvement of $\sim$ 3% compared to SciBERT on classifying an article subject based on its abstract. Differently, MaterialsBERT167 fine-tuned PubMedBERT161 on 2.4M abstracts extracted from material science articles using a masked training approach. MaterialsBERT was then used to generate a contextual encoded representation of the abstract to train a named entity relationship (NER) model. Reporting an increase in precision on their test dataset of $\sim$ 1%, $\sim$ 2%, and $\sim$ 10% for PubMedBERT, MatBERT, and ChemBERT, respectively.
@@
Considerable effort was also employed to develop such models focused on biology tasks. They follow a similar trend of training an LLM model on a large training corpus like Wikipedia, scientific data extracted from a scientific database, and textbooks, followed by fine-tuning specific downstream tasks. Shin et al. 170 pretrained different Megatron-LM318, another BERT-like LLM, models with different sizes, originating the BioMegatron family of models. Considering models with 345M, 800M, and 1.2B parameters and vocabularies with 30k and 50k tokens, they compared the influence of model size on its performance. These models were pretrained using abstracts from the PubMed dataset and full-text scientific articles from PubMed Central (PMC), similar to BioBERT171. The largest 1.2B training started from a BERT-uncased checkpoint to save time. Surprisingly, the largest model did not perform better than the smaller ones. In fact, the 345M parameters using the 50k tokens vocabulary consistently outperformed other models in NER and relation extraction (RE) tasks. Interestingly, the model size seems relevant for the SQuAD319 dataset, showing that LLMs trained on small, domain-specific datasets may lack generalization abilities. BioBERT171 used data from Wikipedia, textbooks, PubMed abstracts, and PMC full-text corpus to pretrain a BERT architecture. Outperforming the original BERT in all benchmarks tested and even SOTA models in some benchmarks (NCBI disease, 2010 i2b2/VA, BC5CDR, BC4CHEMD, BC2GM, JNLPBA, LINNAEUS, Species-800). Peng et al. 320 developed a multi-task BERT evaluated on the BLUE163 benchmark called BlueBERT. The model was pretrained on PubMed abstracts and MIMIC-III313 and fine-tuned on BLUE tasks, achieving similar results as BioBERT in multiple benchmarks. PubMedBERT161, following the approach adopted in SciBERT, developed a domain-specific vocabulary. They used 14M abstracts extracted from PubChem to pretrain PubMedBERT. In addition, they curated and grouped multiple biomedical datasets to develop BLURB, their comprehensive benchmark for biomedical NLP tasks, such as NER, sentence similarity, document classification, and question-answering. Gu et al. 161 showed that PubMedChem considerably outperformed other LLMs in the BLURB benchmark, mainly in PubMedQA and BioQSA datasets. The second best-evaluated model in these datasets was BioBERT, evidencing the importance of domain-specific knowledge to high-performance LLMs.
@@
The models discussed previously in the section use an encoder-only LLM to encode information from scientific data. The learned encoded representation is then further used to make predictions by feeding it to simpler models. A different approach involves training a decoder model. Using a decoder model aims to address the lack of generation ability of the encoder-only models. By implementing a decoder, the model can address a different group of generative tasks, such as question-answering and document classification, using generated labels instead of following a pre-defined set of possible classes. Darwin153 aims to create domain-specific LLMs for natural science by first fine-tuning LLaMA-7B on FAIR, a general Question-Answering (QA) dataset, and then fine-tuning it on scientific QA datasets. Scientific QA instructions were obtained from SciQ321 and generated by the authors using the Scientific Instruction Generation (SIG) model also developed by the authors. Their SIG model is fine-tuned from Vicuna-7B to convert full-text scientific papers into a set of QA pairs. This approach presented an improved performance of DARWIN in several regression and classification benchmarks. Interestingly, Using LLaMA-7B fine-tuned on FAIR only – with no scientific-specific QA fine-tuning – led to similar results in six out of nine benchmarks with metrics available for both models.
@@
In biology, BioGPT156 pretrained a GPT-2 model architecture using 15M abstracts from PubChem corpus. BioGPT was evaluated on four tasks across five benchmarks: end-to-end relation extraction on BC5CDR, KD-DTI, and DDI, question answering on PubMedQA, document classification on HoC, and text generation on the previously listed benchmarks. fine-tuning BioGPT on downstream tasks (except text generation) showed its superiority against other encoder-only LLMs in the biomedical domain, such as BioBERT and PubMedBERT. Focusing on text generation, the authors qualitatively compared BioGPT and GPT-2 generations. They conclude that BioGPT demonstrates better text generation. However, no metric was used, making the comparison difficult to measure. With a similar idea, Wu et al. 154 first pretrained an LLaMA2 model with the MedC-k dataset, consisting of 4.8M academic papers and 30K textbooks. In the sequence, the model was aligned through instruction tuning using the MedC-I dataset, a set of medical QA problems. PMC-LLaMA154 outperforms LLaMa-2 and ChatGPT in multiple biomedical QA benchmarks despite its size $\sim$ 10 times smaller. Interestingly, the authors show that the performance of PMC-LLaMA on MedQA322, MedMCQA323, and PubMedQA324 benchmarks progressively increases as more knowledge is injected, the model size is increased, and more instructions are included in instruction tuning.
@@ 3.7 The use of ChatGPT in Chemistry
With the rise of ChatGPT, we review here how many researchers have wanted to test the capability of such an accessible decoder-only LLM. Castro Nascimento and Pimentel 325 wrote the first notable paper on ChatGPT’s impact on Chemistry. The authors emphasize that LLMs, trained on extensive, uncurated datasets potentially containing errors or secondary sources, may include inaccuracies limiting their ability to predict chemical properties or trends. The paper highlighted that while LLMs could generate seemingly valid responses, they lacked true reasoning or comprehension abilities and would perpetuate existing errors from their training data. However, the authors suggested that these limitations could be addressed in the future. The work serves as a benchmark to qualitatively assess improvements in generative pretrained transformers. For example, five tasks were given to ChatGPT (GPT-3). The accuracy for converting compound names to SMILES representations and vice versa was about 27%, with issues in differentiating alkanes and alkenes, benzene and cyclohexene, or cis and trans isomers. ChatGPT found reasonable octanol-water partition coefficients with a 31% mean relative error, and a 58% hit rate for coordination compounds’ structural information. It had a 100% hit rate for polymer water solubility and a 60% hit rate for molecular point groups. Understandably, the best accuracies were achieved with widely recognized topics. The authors concluded that neither experimental nor computational chemists should fear the development of LLMs, nor task automation. Instead, they encouraged the improvement of dedicated AI tools for specific problems, with their integration into research, as facilitators.
@@
The sole use of ChatGPT for chemistry and any apparent advancement in chemistry seems somewhat limited. Humphry and Fuller 326, Emenike and Emenike 327, and Fergus et al. 328 address the use of ChatGPT in chemical education. There are also a few papers that describe the use of ChatGPT on a specific area of chemical research,329, 330, 331 where fine-tuned GPT models are used to direct the synthesis and functional optimization of Metal Organic Framewoks (MOFs), merging computational modeling with empirical chemistry research. Deb et al. 332 provide a thorough assessment of ChatGPT’s effectiveness in material science through various tasks in computational material science, though the perspective is subjective. ChatGPT assisted in computational materials science by providing crystal space groups for simple compounds, suggesting inputs for simulations, refining analysis with specific rules, and identifying resources. The authors highlight ChatGPT’s potential to write code for enhancing processes and as a tool for those without prior expertise, particularly in catalyst development for CO2 capture. We highlight three points about using ChatGPT alone. First, reliable outputs require detailed and accurate input, as Deb et al. 332 showed, where ChatGPT struggled to mine or predict crystal structures. Second, standardized methods for reproducing and evaluating GPT work are still unclear. Third, this work naturally leads to the inference that additional chemical tools or agents are needed to achieve more complex thinking, as described by Bloom’s Taxonomy.333, 334 Bloom’s taxonomy categorizes educational goals into hierarchical levels: Remembering, Understanding, Applying, Analyzing, Evaluating, and Creating. Remembering involves recalling facts, Understanding means restating a concept, and Applying uses a theory in a new situation. Analyzing breaks down a mechanism into parts to understand connections and differences. Evaluating involves making value judgments, and Creating entails building new structures from diverse elements. Currently, LLMs and autonomous agents are limited in replicating higher-level thinking compared to human understanding. To better quantify their capabilities in this area, we therefore recommend adopting a quality metric based on Bloom’s taxonomy. This educational framework can serve as an effective tool for evaluating the sophistication of LLMs and autonomous agents, especially when tackling complex chemical challenges. Bloom’s taxonomy would thereby provide a structured approach to assess higher-order thinking and reasoning.
@@ 3.7.1 Automation
The evolution of artificial intelligence in chemistry has fueled the potential for automating scientific processes. For example, in 2019, Coley et al. 335 developed a flow-based synthesis robot proposing synthetic routes and assembling flow reaction systems, tested on medically relevant molecules, and in 2020, Gromski et al. 336 provided a useful exploration on how chemical robots could outperform humans when executing chemical reactions and analyses. They developed the Chemputer, a programmable batch synthesis robot handling reactions like peptide synthesis and Suzuki coupling. In 2021, Grisoni et al. 337 combined deep learning-based molecular generation with on-chip synthesis and testing. The Automated Chemical Design (ACD) framework by Goldman et al. 338 provides a useful taxonomy for automation and experimental integration levels. Thus, automation promises to enhance productivity through increased efficiency, error reduction, and the ability to handle complex problems, as described in several excellent reviews regarding automation in chemistry,339, 340, 341, 342, 343, 344, 345
This increased productivity may be the only possible approach to exploring the vastness of all chemical space. To fully leverage AI in property prediction, inverse design, and synthesis prediction, it must be integrated with automated synthesis, purification, and testing. This automation should be high-throughput and driven by AI-based autonomous decision-making (sometimes called “lights-out” automation). Janet et al. 340 highlighted challenges in multi-step reactions with intermediate purifications, quantifying uncertainty, and the need for standardized recipe formats. They also stated the limitations of automated decision-making. Organa346 addresses some of these challenges. It can significantly reduce physical workload and improve users’ lab experience by automating diverse common lab routine tasks such as solubility assessment, pH measurement, and recrystallization. Organa interacts with the user through text and audio. The commands are converted into a detailed LLM prompt and used to map the goal to the robot’s instructions. Interestingly, Organa is also capable of reasoning over the instructions, giving feedback about the experiments, and producing a written report with the results.

### v2

@@ 3.6 Textual scientific LLMs
LLMs are large deep neural network models. Therefore, their superior performance against other smaller models in classical machine learning tasks is somewhat expected. Addressing tasks that seemed impossible to tackle in the past better illustrates LLM’s outstanding capabilities. That is, despite accurately predicting properties given well-structured information such as molecular features and descriptors, LLMs can access less structured data and extract information from simple natural language. For instance, instead of developing a restricted representation, Huang et al. 138 used clinical annotations from the MIMIC-III336 dataset to predict patient readmission. ClinicalBERT implements a pretrained BERT architecture using a combination of masked language modeling and next-sentence prediction training schemes followed by supervised fine-tuning on readmission prediction. Or, instead of pretraining, Zhao et al. 121 fine-tuned LLaMA on epilepsy data using an instruction-following approach to create EpilepsyLLM. Similarly, SciBERT134 and ScholarBERT137 adapted BERT to scientific literature. Beltagy et al. 134 created a tokenizer specific for scientific texts extracted from scientific articles from semantic scholar180 and showed that SciBERT outperforms fine-tuned BERT-base105 in every task investigated. It shows that vocabulary quality plays an important role in the model’s performance. A few years later, Hong et al. 137 pretrained BERT using RoBERTa337 optimizations to improve pretraining performance. ScholarBERT was pretrained on scientific articles from Public.Resource.Org, Inc and further fine-tuned on the tasks used for evaluation. Despite using a much larger dataset, ScholarBERT did not outperform other LLMs trained with narrower domain datasets. Even against SciBERT, another LLM trained on multi-domain articles, ScholarBERT could not perform relatively better in tasks from biomedical, computer science, and material domains. However, ScholarBERT showed significant improvement in the ScienceExamCER338 dataset, which contains semantic class annotations in a training corpus of 3rd to 9th grade science exam questions for named entity recognition (NER) tasks.
@@
On the other hand, instead of pretraining LLMs on a large domain scope, models focused on only one field were also developed. Guo et al. 141 argue that manually curating structured datasets is a sub-optimal, time-consuming, and labor-intensive task. Therefore, they automated data extraction and annotation from scientific papers using ChemDataExtractor339 and their in-house annotation tool340. ChemBERT141 was pretrained using a BERT model to encode chemical reaction information, followed by fine-tuning a NER head. ChemBERT outperformed other models such as BERT105 and BioBERT143 in the product extraction task, presenting an improvement of $\sim$ 6% in precision. For the product role labeling task – given the extracted compound, which role does it play in the reaction? –, it increases precision in $\sim$ 5% when compared to the same models. The increase in performance might be due to training in a narrower domain, which allows the model to learn specific traits from the data instead of being more generalist. MatSciBERT132 and MaterialsBERT139 followed a similar strategy, and BatteryBERT,197 similarly outperformed original BERT models with application to specific battery tasks. Gupta et al. 132 developed MatSciBERT by fine-tuning SciBERT134 on the Material Science Corpus (MSC), their curated dataset focused on materials extracted from Elsevier’s scientific papers. MatSciBERT shows an accuracy improvement of $\sim$ 3% compared to SciBERT on classifying an article subject based on its abstract. Differently, MaterialsBERT139 fine-tuned PubMedBERT133 on 2.4M abstracts extracted from material science articles using a masked training approach. MaterialsBERT was then used to generate a contextual encoded representation of the abstract to train a named entity relationship (NER) model. Reporting an increase in precision on their test dataset of $\sim$ 1%, $\sim$ 2%, and $\sim$ 10% for PubMedBERT, MatBERT, and ChemBERT, respectively.
@@
Considerable effort was also employed to develop such models focused on biology tasks. They follow a similar trend of training an LLM model on a large training corpus like Wikipedia, scientific data extracted from a scientific database, and textbooks, followed by fine-tuning specific downstream tasks. Shin et al. 142 pretrained different Megatron-LM341, another BERT-like LLM, models with different sizes, originating the BioMegatron family of models. Considering models with 345M, 800M, and 1.2B parameters and vocabularies with 30k and 50k tokens, they compared the influence of model size on its performance. These models were pretrained using abstracts from the PubMed dataset and full-text scientific articles from PubMed Central (PMC), similar to BioBERT143. The largest 1.2B training started from a BERT-uncased checkpoint to save time. Surprisingly, the largest model did not perform better than the smaller ones. In fact, the 345M parameters using the 50k tokens vocabulary consistently outperformed other models in NER and relation extraction (RE) tasks. Interestingly, the model size seems relevant for the SQuAD342 dataset, showing that LLMs trained on small, domain-specific datasets may lack generalization abilities. BioBERT143 used data from Wikipedia, textbooks, PubMed abstracts, and PMC full-text corpus to pretrain a BERT architecture. Outperforming the original BERT in all benchmarks tested and even SOTA models in some benchmarks (NCBI disease, 2010 i2b2/VA, BC5CDR, BC4CHEMD, BC2GM, JNLPBA, LINNAEUS, Species-800). Peng et al. 343 developed a multi-task BERT evaluated on the Biomedical Language Understanding Evaluation (BLUE)135 benchmark called BlueBERT. The model was pretrained on PubMed abstracts and MIMIC-III336 and fine-tuned on BLUE tasks, achieving similar results as BioBERT in multiple benchmarks. PubMedBERT133, following the approach adopted in SciBERT, developed a domain-specific vocabulary. They used 14M abstracts extracted from PubChem to pretrain PubMedBERT. In addition, they curated and grouped multiple biomedical datasets to develop BLURB, their comprehensive benchmark for biomedical NLP tasks, such as NER, sentence similarity, document classification, and question-answering. Gu et al. 133 showed that PubMedChem considerably outperformed other LLMs in the BLURB benchmark, mainly in PubMedQA and BioQSA datasets. The second best-evaluated model in these datasets was BioBERT, evidencing the importance of domain-specific knowledge to high-performance LLMs.
@@
The models discussed previously in the section use an encoder-only LLM to encode information from scientific data. The learned encoded representation is then further used to make predictions by feeding it to simpler models. A different approach involves training a decoder model. Using a decoder model aims to address the lack of generation ability of the encoder-only models. By implementing a decoder, the model can address a different group of generative tasks, such as question-answering and document classification, using generated labels instead of following a pre-defined set of possible classes. Darwin124 aims to create domain-specific LLMs for natural science by first fine-tuning LLaMA-7B on FAIR, a general Question-Answering (QA) dataset, and then fine-tuning it on scientific QA datasets. Scientific QA instructions were obtained from SciQ344 and generated by the authors using the Scientific Instruction Generation (SIG) model also developed by the authors. Their SIG model is fine-tuned from Vicuna-7B to convert full-text scientific papers into a set of QA pairs. This approach presented an improved performance of DARWIN in several regression and classification benchmarks. Interestingly, Using LLaMA-7B fine-tuned on FAIR only – with no scientific-specific QA fine-tuning – led to similar results in six out of nine benchmarks with metrics available for both models.
@@
In biology, BioGPT127 pretrained a GPT-2 model architecture using 15M abstracts from PubChem corpus. BioGPT was evaluated on four tasks across five benchmarks: end-to-end relation extraction on BC5CDR, KD-DTI, and DDI, question answering on PubMedQA, document classification on HoC, and text generation on the previously listed benchmarks. fine-tuning BioGPT on downstream tasks (except text generation) showed its superiority against other encoder-only LLMs in the biomedical domain, such as BioBERT and PubMedBERT. Focusing on text generation, the authors qualitatively compared BioGPT and GPT-2 generations. They conclude that BioGPT demonstrates better text generation. However, no metric was used, making the comparison difficult to measure. With a similar idea, Wu et al. 125 first pretrained an LLaMA2 model with the MedC-k dataset, consisting of 4.8M academic papers and 30K textbooks. In the sequence, the model was aligned through instruction tuning using the MedC-I dataset, a set of medical QA problems. PMC-LLaMA125 outperforms LLaMa-2 and ChatGPT in multiple biomedical QA benchmarks despite its size $\sim$ 10 times smaller. Interestingly, the authors show that the performance of PMC-LLaMA on MedQA345, MedMCQA346, and PubMedQA347 benchmarks progressively increases as more knowledge is injected, the model size is increased, and more instructions are included in instruction tuning.
@@ 3.7 The use of ChatGPT in Chemistry
With the rise of ChatGPT, we review here how many researchers have wanted to test the capability of such an accessible decoder-only LLM. Castro Nascimento and Pimentel 348 wrote the first notable paper on ChatGPT’s impact on Chemistry. The authors emphasize that LLMs, trained on extensive, uncurated datasets potentially containing errors or secondary sources, may include inaccuracies limiting their ability to predict chemical properties or trends. The paper highlighted that while LLMs could generate seemingly valid responses, they lacked true reasoning or comprehension abilities and would perpetuate existing errors from their training data. However, the authors suggested that these limitations could be addressed in the future. The work serves as a benchmark to qualitatively assess improvements in generative pretrained transformers. For example, five tasks were given to ChatGPT (GPT-3). The accuracy for converting compound names to SMILES representations and vice versa was about 27%, with issues in differentiating alkanes and alkenes, benzene and cyclohexene, or cis and trans isomers. ChatGPT found reasonable octanol-water partition coefficients with a 31% mean relative error, and a 58% hit rate for coordination compounds’ structural information. It had a 100% hit rate for polymer water solubility and a 60% hit rate for molecular point groups. Understandably, the best accuracies were achieved with widely recognized topics. The authors concluded that neither experimental nor computational chemists should fear the development of LLMs, nor task automation. Instead, they encouraged the improvement of dedicated AI tools for specific problems, with their integration into research, as facilitators.
@@
The sole use of ChatGPT for chemistry and any apparent advancement in chemistry seems somewhat limited. Humphry and Fuller 349, Emenike and Emenike 350, and Fergus et al. 351 address the use of ChatGPT in chemical education. There are also a few papers that describe the use of ChatGPT on a specific area of chemical research, namely with the synthesis and functional optimization of Metal Organic Framewoks (MOFs), merging computational modeling with empirical chemistry research.352, 353 There is also some overlap with fine-tuned GPT models,354 within this field.355 Deb et al. 356 provide a thorough assessment of ChatGPT’s effectiveness in material science through various tasks in computational material science, though the perspective is subjective. ChatGPT assisted in computational materials science by providing crystal space groups for simple compounds, suggesting inputs for simulations, refining analysis with specific rules, and identifying resources. The authors highlight ChatGPT’s potential to write code for enhancing processes and as a tool for those without prior expertise, particularly in catalyst development for CO2 capture. We highlight three points about using ChatGPT alone. First, reliable outputs require detailed and accurate input, as Deb et al. 356 showed, where ChatGPT struggled to mine or predict crystal structures. Second, standardized methods for reproducing and evaluating GPT work are still unclear. Third, this work naturally leads to the inference that additional chemical tools or agents are needed to achieve more complex thinking, as described by Bloom’s Taxonomy.357, 358 Bloom’s taxonomy categorizes educational goals into hierarchical levels: Remembering, Understanding, Applying, Analyzing, Evaluating, and Creating. Remembering involves recalling facts, Understanding means restating a concept, and Applying uses a theory in a new situation. Analyzing breaks down a mechanism into parts to understand connections and differences. Evaluating involves making value judgments, and Creating entails building new structures from diverse elements. Currently, LLMs and autonomous agents are limited in replicating higher-level thinking compared to human understanding. To better quantify their capabilities in this area, we therefore recommend adopting a quality metric based on Bloom’s taxonomy. This educational framework can serve as an effective tool for evaluating the sophistication of LLMs and autonomous agents, especially when tackling complex chemical challenges. Bloom’s taxonomy would thereby provide a structured approach to assess higher-order thinking and reasoning.
@@ 3.7.1 Automation
The evolution of artificial intelligence in chemistry has fueled the potential for automating scientific processes. For example, in 2019, Coley et al. 359 developed a flow-based synthesis robot proposing synthetic routes and assembling flow reaction systems, tested on medically relevant molecules, and in 2020, Gromski et al. 360 provided a useful exploration on how chemical robots could outperform humans when executing chemical reactions and analyses. They developed the Chemputer, a programmable batch synthesis robot handling reactions like peptide synthesis and Suzuki coupling. In 2021, Grisoni et al. 361 combined deep learning-based molecular generation with on-chip synthesis and testing. The Automated Chemical Design (ACD) framework by Goldman et al. 362 provides a useful taxonomy for automation and experimental integration levels. Thus, automation promises to enhance productivity through increased efficiency, error reduction, and the ability to handle complex problems, as described in several excellent reviews regarding automation in chemistry,363, 364, 365, 366, 367, 368, 369
This increased productivity may be the only possible approach to exploring the vastness of all chemical space. To fully leverage AI in property prediction, inverse design, and synthesis prediction, it must be integrated with automated synthesis, purification, and testing. This automation should be high-throughput and driven by AI-based autonomous decision-making (sometimes called “lights-out” automation). Janet et al. 364 highlighted challenges in multi-step reactions with intermediate purifications, quantifying uncertainty, and the need for standardized recipe formats. They also stated the limitations of automated decision-making. Organa370 addresses some of these challenges. It can significantly reduce physical workload and improve users’ lab experience by automating diverse common lab routine tasks such as solubility assessment, pH measurement, and recrystallization. Organa interacts with the user through text and audio. The commands are converted into a detailed LLM prompt and used to map the goal to the robot’s instructions. Interestingly, Organa is also capable of reasoning over the instructions, giving feedback about the experiments, and producing a written report with the results.